### v3

@@ Appendix B Training of LSTM
The input to the LSTM is the state and the action chosen at time $t$ and it outputs the corresponding Q-values. We made a custom one-hot encoding of the action to nearly match the dimension of the state vector. For example, in the Highway environment, the state vector is 26 dimensional and there are 5 actions. We encode the selected action to a 25-dimensional vector with one filling in the corresponding 5 positions of the action and the rest being zero. We then condition the state vector with the action vector and train the LSTM. An overview of this dataflow is outlined in Figure 4.
@@
Table 7: Table showing the LSTM Optimizer details
@@
Table 8: Table showing the VAE Optimizer details
@@
Table 9: VAE encoder architecture
@@
Table 10: VAE decoder architecture
@@
Experiments were conducted across three distinct environments, each employing different modes of personalisation. For the 2D Navigation and Highway environments, we implemented experiments in avoidance, preference, and mixed modes. However, for the Pong environment, personalisation was conducted in preference mode. The LSTM loss function for all modes across the 2D Navigation, Highway, and Pong environments is depicted in Figures 5, 6, and 7, respectively.
@@ Appendix C Training of VAE
VAE is pretrained and only used if the observation of the environment is pixel-based. Training data is the same as the training data of the task policy. We use the conventional VAE loss functions which consist of reconstruction loss and KL-divergence loss. The encoder and decoder architectures are tabulated in Tables 9, 10 respectively. A fully connected linear layer is used to map the output from the encoder to a dimension of 512 subsequently the output is mapped to a latent dimension (set as 16) using another linear layer. Optimiser details of the VAE is shown in Table 8. The loss curve is shown in Figure 8.
@@ D.4 Entropy Weighted
The term ×$\mathcal{H}^{*}\times\pi_{\phi}(a|s_{t})$ represents the weighted contribution of the task policy, and $(1-\mathcal{H}^{*})\times\pi_{\psi}(a|s_{t})$ represents the weighted contribution of the Intent-specific policy.
Appendix E Analysis of other choices of policy fusion
@@
$KL\left(\pi_{\phi}(a|s)\middle\|\pi_{f}(a|s)\right)\leq\log\left(Z\right)+\left(\dfrac{Q^{*}\delta+\epsilon T_{\phi}}{T_{\phi}T_{\psi}}\right)+\log\left(\zeta\right)\ \forall a\in\mathcal{A},s\in\mathcal{S}$,
@@
$=\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\dfrac{QT_{\psi}-Q^{\prime}T_{\phi}+QT_{\phi}-QT_{\phi}}{T_{\phi}T_{\psi}}\right)+\log\left(\zeta\right)$
$\leq\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\dfrac{Q\delta+\epsilon T_{\phi}}{T_{\phi}T_{\psi}}\right)+\log\left(\zeta\right)$
@@
Appendix F MORL baseline Implementation Details
To implement the MORL baseline, we construct the reward as a convex sum of the rewards of both objectives as $r=\alpha r_{e}+(1-\alpha)r_{h}$, where $r_{e}$ is the reward obtained from the environment and the $r_{h}$ is the reward obtained from the LSTM as in Equation (1). To match with our zero-shot approach, we change the reward of the transitions of the trajectories saved for the training of LSTM. We then populated the new transitions in the replay buffer and trained a new DQN. We choose $\alpha=0.5$ in all our experiments to balance both objective and normalised the human reward components between $[-1,1]$.
Appendix G Additional experiments
In this section, we summarise some additional experiments conducted. Firstly, we checked the behaviour of the static fusion when the LSTM temperature ($T_{\psi}$) is kept as $T_{min}$ in 2D Navigation environment. In this experiment, we found that the fused policy is dominated by human objectives as shown in Table 12. This is because a low temperature results in a sharper Boltzmann probability distribution. Therefore the intent-policy dominates over the other.
Secondly, We investigated the behaviour change with different reward component weighting ($\alpha$) in the MORL baseline. Since MORL, tries to optimise different objectives simultaneously, the weighting of the reward component plays a significant role. If the weighting of the reward component is high, the objective corresponding to that reward is highly prioritised. We empirically tested this in the Highway environment with weightings as $\alpha=0.3,0.5$ and $0.7$. As the value of the $\alpha$ increases, the policy obtained from MORL shifts more towards the human objective as in Table 13.
Appendix H DQN Training
DQN was utilised to obtain the task-specific policy. For the 2D navigation environment, which involves image-based observations, we adopted the CNN network as detailed in Table 11. In the case of the Pong and Highway environments, an MLP was implemented, featuring three hidden layers with sizes of 256, 512, and 512. We use the ReLU activation function after each layer. We set the minimum epsilon at 0.10, with a decay rate of 0.995. The reward curve for each environment is depicted in Figure 9.
Table 11: Table showing the CNN architecture of DQN
Particulars Value
Channel 32,64,64
Kernel size 11,5,3
Stride 2,2,1
Padding 0,0,0
FC hidden layer 512
(a) 2D Navigation
(b) Highway
(c) Pong
Figure 9: Smoothed reward curve of DQN for 2D Navigation, Highway and Pong
Table 12: Result of static fusion when $T_{\psi}=T_{min}$ in 2D Navigation. The up arrow $\uparrow$ indicates higher is better, and the down arrow $\downarrow$ indicates lower is better. The result averaged over 10 seeds with each seed containing 300 episodes with $T_{\phi}=0.4$, $T_{min}=1$, $T_{max}=10$, and $\eta=0$
Mode Method Desired region $\uparrow$ Undesired region $\downarrow$ Score $\uparrow$
@@ Appendix H DQN Training
Mixed Static $2.372\pm 0.123$ $000.000\pm 0$ 0.000 $0.739\pm 0.014$
Avoidance Static - $000.000\pm 0$ 0.000 $0.974\pm 0.011$
Table 13: Results correspond to the Highway environment. The up arrow $\uparrow$ indicates higher values are better and the down arrow $\downarrow$ indicates lower values are better. The hyper-parameters chosen are $T_{\phi}=0.6,T_{max}=5,T_{min}=0.3,\eta=0$
Mode $\alpha$ Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$
Avoidance 0.3 - $16.25\pm 3.45$ $0.15\pm 0.03$ $37.95\pm 0.68$
0.5 - $5.80\pm 1.07$ $0.14\pm 0.23$ $37.44\pm 0.52$
0.7 - $2.75\pm 0.49$ $0.47\pm 0.02$ $33.90\pm 0.91$

### v4

@@ Appendix B Training of LSTM
The input to the LSTM is the state and the action chosen at time $t$ and it outputs the corresponding Q-values. We made a custom one-hot encoding of the action to nearly match the dimension of the state vector. For example, in the Highway environment, the state vector is 26 dimensional and there are 5 actions. We encode the selected action to a 25-dimensional vector with one filling in the corresponding 5 positions of the action and the rest being zero. We then condition the state vector with the action vector and train the LSTM. An overview of this dataflow is outlined in Fig 4.
@@
Table 8: LSTM optimizer details.
@@
Table 9: VAE optimizer details.
@@
Table 10: VAE encoder architecture.
@@
Table 11: VAE decoder architecture.
@@
Experiments were conducted across three distinct environments, each employing different modes of personalisation. For the 2D Navigation and Highway environments, we implemented experiments in avoidance, preference, and mixed modes. However, for the Pong environment, personalisation was conducted in preference mode. The LSTM loss function for all modes across the 2D Navigation, Highway, and Pong environments is depicted in Fig 5, 6, and 7, respectively.
@@ Appendix C Training of VAE
VAE is pretrained and only used if the observation of the environment is pixel-based. Training data is the same as the training data of the task policy. We use the conventional VAE loss functions which consist of reconstruction loss and KL-divergence loss. The encoder and decoder architectures are tabulated in Tables 10, 11 respectively. A fully connected linear layer is used to map the output from the encoder to a dimension of 512 subsequently the output is mapped to a latent dimension (set as 16) using another linear layer. Optimiser details of the VAE is shown in Table 9. The loss curve is shown in Fig 8.
@@ D.4 Entropy Weighted
The term $\mathcal{H}^{*}\times\pi_{\phi}(a|s_{t})$ represents the weighted contribution of the task policy, and $(1-\mathcal{H}^{*})\times\pi_{\psi}(a|s_{t})$ represents the weighted contribution of the Intent-specific policy.
Appendix E Proof of main theorem and lemma
Theorem (1).
Proof.
$+\dfrac{1}{2}\log\left(\zeta\right)$
Lemma (2).
Let $\pi_{\psi}$ and $\pi_{\phi}$ be intent-specific and task-specific policies respectively, and assume $\pi_{\psi}$ is not a random policy. Then $\pi_{\phi}$ and $\tfrac{1}{Z}\pi_{\psi}\pi_{\phi}$, where $Z$ is a normalising factor, are not invariant policies under any condition.
Proof.
We compute the KL divergence between $\pi_{\phi}$ and $\tfrac{1}{Z}\pi_{\psi}\pi_{\phi}$:
$KL\!\left(\pi_{\phi}\;\middle\|\;\tfrac{\pi_{\phi}\pi_{\psi}}{Z}\right)$ $=\sum_{a}\pi_{\phi}(a)\log\left(\frac{\pi_{\phi}(a)}{\tfrac{1}{Z}\pi_{\phi}(a)\pi_{\psi}(a)}\right)$
$=\sum_{a}\pi_{\phi}(a)\log\left(\frac{Z}{\pi_{\psi}(a)}\right)$
$=\log(Z)-\sum_{a}\pi_{\phi}(a)\log\pi_{\psi}(a),$
where the normalising factor is
$Z=\sum_{a}\pi_{\phi}(a)\pi_{\psi}(a).$
For $\pi_{\phi}$ and $\tfrac{1}{Z}\pi_{\psi}\pi_{\phi}$ to be invariant, we require
$KL\!\left(\pi_{\phi}\;\middle\|\;\tfrac{\pi_{\phi}\pi_{\psi}}{Z}\right)=0.$
That is,
$\log\left(Z\right)=\sum_{a}\pi_{\phi}(a)\log\pi_{\psi}(a).$
Equivalently,
$\log\!\left(\sum_{a}\pi_{\phi}(a)\pi_{\psi}(a)\right)=\sum_{a}\pi_{\phi}(a)\log\pi_{\psi}(a),$
which can be rewritten as
$\log\Big(\mathbb{E}_{\pi_{\phi}}[\pi_{\psi}]\Big)=\mathbb{E}_{\pi_{\phi}}[\log\pi_{\psi}].$
By Jensen’s inequality, since $\log$ is concave, we have
$\mathbb{E}_{\pi_{\phi}}[\log\pi_{\psi}]\;\leq\;\log\Big(\mathbb{E}_{\pi_{\phi}}[\pi_{\psi}]\Big),$
with equality if and only if $\pi_{\psi}(a)$ is constant almost surely under $\pi_{\phi}$, i.e., if $\pi_{\psi}$ is a random (uniform) policy. This contradicts the assumption that $\pi_{\psi}$ is not a random policy.
Therefore, the KL divergence cannot be zero, and hence $\pi_{\phi}$ and $\tfrac{1}{Z}\pi_{\psi}\pi_{\phi}$ are not invariant policies. ∎
Appendix F Analysis of other choices of policy fusion
@@
$KL\left(\pi_{\phi}(a|s)\middle\|\tfrac{\pi_{\phi}(a|s)\pi_{\psi}(a|s)}{Z}\right)\leq\log\left(Z\right)+\left(\dfrac{Q^{*}\delta+\epsilon T_{\phi}}{T_{\phi}T_{\psi}}\right)+\log\left(\zeta\right)\ \forall a\in\mathcal{A},s\in\mathcal{S}$,
@@
$=\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\dfrac{QT_{\psi}-Q^{\prime}T_{\phi}+QT_{\phi}-QT_{\phi}}{T_{\phi}T_{\psi}}\right)$
$\qquad+\log\left(\zeta\right)$
$\leq\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\dfrac{Q\delta+\epsilon T_{\phi}}{T_{\phi}T_{\psi}}\right)$
$\qquad+\log\left(\zeta\right)$
@@
Appendix G MORL baseline Implementation Details
To implement the MORL baseline, we construct the reward as a convex sum of the rewards of both objectives as $r=\alpha r_{e}+(1-\alpha)r_{h}$, where $r_{e}$ is the reward obtained from the environment and the $r_{h}$ is the reward obtained from the LSTM as in Equation (1). To match with our zero-shot approach, we change the reward of the transitions of the trajectories saved for the training of LSTM. We then populated the new transitions in the replay buffer and trained a new DQN. We choose $\alpha=0.5$ in all our experiments to balance both objective and normalised the human reward components between [1,1][-1,1].
Table 12: Results of static fusion when $T_{\psi}=T_{\min}$ in 2D Navigation. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better. Results averaged over 10 seeds, each with 300 episodes. Hyperparameters: $T_{\phi}=0.4$, $T_{\min}=1$, $T_{\max}=10$, $\eta=0$.
Mode Method
Desired region
visits $\uparrow$
Undesired region
visits $\downarrow$
Score
$\uparrow$
@@ Appendix G MORL baseline Implementation Details
Mixed Static $2.372\pm 0.123$ $0.000\pm 0.000$ $0.739\pm 0.014$
Avoidance Static $0.000\pm 0.000$ $0.974\pm 0.011$
Table 13: Results in the Highway environment with different values of $\alpha$. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better. Hyperparameters: $T_{\phi}=0.6$, $T_{\max}=5$, $T_{\min}=0.3$, $\eta=0$.
Mode $\alpha$
Desired lane
visits $\uparrow$
Undesired lane
visits $\downarrow$
Hits
$\downarrow$
Score
$\uparrow$
Avoidance 0.3 $16.25\pm 3.45$ $0.15\pm 0.03$ $37.95\pm 0.68$
0.5 $5.80\pm 1.07$ $0.14\pm 0.23$ $37.44\pm 0.52$
0.7 $2.75\pm 0.49$ $0.47\pm 0.02$ $33.90\pm 0.91$
@@
Appendix H Additional experiments
In this section, we report a set of supplementary experiments designed to provide further insights into the behaviour of our method and baselines.
Firstly, we examined the effect of fixing the LSTM temperature ($T_{\psi}$) at its minimum value $T_{\min}$ in the 2D Navigation environment. As shown in Table 12, the fused policy becomes dominated by the human objective. This is expected, since a low temperature sharpens the Boltzmann distribution, amplifying the influence of the intent-specific policy relative to the task policy.
Secondly, we analysed the sensitivity of the MORL baseline to different reward weightings ($\alpha$). Because MORL optimises multiple objectives simultaneously, the choice of $\alpha$ strongly affects the trade-off between task and human rewards. In the Highway environment, we tested $\alpha\in\{0.3,0.5,0.7\}$. As reported in Table 13, increasing $\alpha$ shifts the policy more towards the human objective, confirming the intuitive role of weighting in MORL.
Effect of $\eta$: Table 14 presents results for different values of $\eta$ under dynamic fusion. Theoretically, a higher $\eta$ increases the likelihood that $T_{\psi}$ remains low (see Fig 2), making the intent-specific policy sharper and more dominant during fusion. The empirical results show a trend consistent with this: in the Preference mode, higher $\eta$ is associated with more visits to the desired lane and fewer to the undesired lane. However, the changes across $\eta$ are relatively modest, and task scores remain broadly similar. In our main experiments we set $\eta=0$, though in other environments an alternative choice may be preferable.
Effect of $T_{\max}$: We also investigated the influence of $T_{\max}$, which controls the maximum temperature in Equation (6). A larger $T_{\max}$ should, in principle, weaken the effect of the intent-specific policy. The results in Table 15 are consistent with this expectation: as $T_{\max}$ increases, the agent shows a reduced tendency to follow the preferred lane and a greater tendency to deviate into undesired lanes. In other words, the level of adherence to human preferences decreases as $T_{\max}$ grows.
Overall, these additional studies confirm the theoretical roles of $T_{\psi}$, $\alpha$, $\eta$, and $T_{\max}$ in shaping behaviour, while also showing that the practical impact of $\eta$ and $T_{\max}$ on performance is limited in the tested settings.
Table 14: Results in the Highway environment with different values of $\eta$. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better for the quantity specified in the column. The hyperparameters are $T_{\phi}=0.6$, $T_{\max}=5$, $T_{\min}=0.3$.
Mode $\eta$
Desired lane
visits $\uparrow$
Undesired lane
visits $\downarrow$
Hits
$\downarrow$
Score
$\uparrow$
Avoidance 0 $28\pm 0.08 0.0$$06\pm 0.01$ $39.83\pm 0.59$
1 $35\pm 0.12 0.0$$08\pm 0.02$ $38.96\pm 0.87$
2 $19\pm 0.11 0.0$$09\pm 0.02$ $38.80\pm 0.51$
Preference 0 $25.14\pm 1.44$ $0.05\pm 0.02$ $40.02\pm 0.61$
1 $24.90\pm 1.52$ $0.10\pm 0.03$ $38.84\pm 0.84$
2 $29.91\pm 1.38$ $0.06\pm 0.01$ $39.27\pm 0.59$
Mixed 0 $21.52\pm 2.00$ $0.85\pm 0.33$ $0.12\pm 0.01$ $38.50\pm 0.43$
1 $23.96\pm 1.67$ $56\pm 0.13 0.0$$07\pm 0.01$ $39.56\pm 0.45$
2 $25.21\pm 2.55$ $27\pm 0.06 0.0$$07\pm 0.04$ $39.15\pm 0.49$
Table 15: Sensitivity analysis of $T_{\max}$ in the Highway environment (Mixed mode). The hyperparameters are $T_{\phi}=0.6$, $T_{\min}=0.3$, and $\eta=0$. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better.
$T_{\max}$
Desired lane
visits $\uparrow$
Undesired lane
visits $\downarrow$
Hits
$\downarrow$
Score
$\uparrow$
10 $19.21\pm 1.42$ $2.02\pm 0.43$ $0.06\pm 0.02$ $40.12\pm 0.59$
15 $17.60\pm 0.90$ $3.42\pm 0.43$ $0.09\pm 0.02$ $39.64\pm 0.45$
25 $9.53\pm 0.82$ $4.65\pm 0.52$ $0.10\pm 0.02$ $39.51\pm 0.53$
35 $14.36\pm 0.65$ $5.32\pm 0.58$ $0.10\pm 0.02$ $40.03\pm 0.59$
Appendix I DQN Training
DQN was utilised to obtain the task-specific policy. For the 2D navigation environment, which involves image-based observations, we adopted the Convolutional Neural Network (CNN) network as detailed in Table 16. In the case of the Pong and Highway environments, an MLP was implemented, featuring three hidden layers with sizes of 256, 512, and 512. We use the ReLU activation function after each layer. We set the minimum epsilon at 0.10, with a decay rate of 0.995.
Table 16: CNN architecture of the DQN agent.
Particulars Value
Channels 32, 64, 64
Kernel sizes 11, 5, 3
Strides 2, 2, 1
Paddings 0, 0, 0
FC hidden layer 512